OpenStudy (anonymous):

SIMPLE QUESTIONS FOR YOU ALL A bus was moving at a speed of 54 km/h. on applying brakes, it stopped in 8 seconds. calculate the speed acquired and distance traveled in this time.

OpenStudy (anonymous):

The speed acquired is 0m/s

OpenStudy (apoorvk):

Speed acquired?????

OpenStudy (anonymous):

Because the bus stopped

OpenStudy (apoorvk):

That be zero, since it stops!

OpenStudy (anonymous):

plzzz do show u r solutions

OpenStudy (anonymous):

distance use, x = v=0.5(u+v)t

OpenStudy (anonymous):

give answer in one comment please its my request

OpenStudy (anonymous):

So x = 0.5* 15 * 8 = 60m

OpenStudy (apoorvk):

Well, you have initial velocity = u = 54 kmph = 54 x (5/18) m/sec = 15 m/sec Final velocity = v = 0 Time = 8 seconds. Use Newton Unkle's equation # 1 ---> \( a = \frac{v-u}{t}\), and fond out the deceleration 'a'. Then, use \(v^2 - u^2 = 2as\) to find out the distance covered, i.e. 's'.

OpenStudy (apoorvk):

(Note of precaution: Newton's equations are only valid for uniform-linear motion, as you may know.)

OpenStudy (king):

sry... use \[x=\frac{v+u}{2} \times t\]

OpenStudy (anonymous):

thanks aproov it proved my solution was wrong

OpenStudy (king):

where v is final speed=0 u is initial speed=15 and t=8

OpenStudy (apoorvk):

Yeah, using the average velocity funda is good actually, @King

OpenStudy (anonymous):

oKAY thanks king to u as well

OpenStudy (apoorvk):

In fact innovative. :]

OpenStudy (king):

yw...thnx @apoorvk

OpenStudy (anonymous):

lol